NicholasHen1615:

Algebra 2: The functions f(x) = x2 – 3 and g(x) = –x2 + 2 are shown on the graph. The graph shows f of x equals x squared minus 3, which is an upward opening parabola with a vertex at 0 comma negative 3 and a point at negative 1 comma negative 2 and a point at 1 comma negative 2. The graph also shows g of x, which is a downward opening parabola with a vertex at 0 comma 2 and a point at negative 1 comma 1 and a point at 1 comma 1. Explain how to modify the graphs of f(x) and g(x) to graph the solution set to the following system of inequalities. How can the solution set be identified? y ≤ x2 – 3 y > –x2 + 2

grezzyy:

The set of inequalities y ≤ x² - 3 and y > -x² + 2 do not have a solution we have: f(x) = x² - 3 g(x) = -x² + 2 Next, we change the equations to inequalities as y ≤ x² - 3 y > -x² + 2 To modify the graph, we then perform the following transformations: Shift the function g(x) down by 2 units Reflect across the x-axis Shift the function g(x) down by 3 units After the modifications in (a), we have: y ≤ x² - 3 and y > -x² + 2 Next, we plot the graph of the inequalities From the graph of the inequalities, the curves of the inequalities have no point of intersection Hence, the set of inequalities do not have a solution

NicholasHen1615:

Alright, Thank you! u up for one more?

grezzyy:

ight

NicholasHen1615:

NicholasHen1615:

Thx

grezzyy:

The f(x) has a degree of 1, hence it is linear in nature while g(x) is a polynomial function because of the third degree. The key features that they have in common are the y-intercept, domain, and range Linear and quadratic equations Linear equations are equations that has a degree of 1 while polynomial equations are equations that has a degree of 2 and above. Given the following functions f(x) = –4x + 5 and; g(x) = x^3 + x^2 – 4x + 5 Since the f(x) has a degree of 1, hence it is linear in nature while g(x) is a polynomial function because of the third degree. The key features that they have in common are the y-intercept, domain and range

NicholasHen1615:

Thank you. And to let you know the first one was wrong :(

grezzyy:

damn

NicholasHen1615:

thx though.

NicholasHen1615:

second one seems right though

NicholasHen1615:

That one was wrong to. :(